OpenStudy (anonymous):

Solve: 8/x^2-1=(3/x+1)+(4/x-1) a. no solution b. 7 c. 1 d. 1/7

hero (hero):

Make sure the denominators are factored first \(\large\frac{8}{(x+1)(x-1)} = \frac{3}{x+1} + \frac{4}{x-1}\)

hero (hero):

Now multiply both sides by x + 1: \(\large\frac{8}{x-1} = 3 + \frac{4(x+1)}{x-1}\)

hero (hero):

Subtract \(\frac{4(x+1)}{x-1}\) from both sides

hero (hero):

\(\large\frac{8}{x-1} - \frac{4x + 4}{x-1} = 3\)

hero (hero):

Combine fractions on the left side: \(\large\frac{4 - 4x}{x-1} = 3\)

hero (hero):

factor the numerator to get: \(\large\frac{-4(x-1)}{x-1} = 3\)

hero (hero):

Cancel the factor of 1 to get \(-4 \ne 3\)

hero (hero):

No solution

hero (hero):

I hope I didn't lose you

OpenStudy (anonymous):

Nope, I got it. thank you very much! that was helpful :)

hero (hero):

@waterineyes is a fan of my methods now.

OpenStudy (anonymous):

Need not to do factorization there..

hero (hero):

It was necessary to continue with my methods

OpenStudy (anonymous):

\[\frac{8}{x^2-1} = \frac{7x + 1}{(x+1)(x-1)} = \frac{7x+1}{x^2 - 1}\] \[x = 1\] But when plugging in 1 back in the equation we get undefined.. So no solution...

hero (hero):

I told you my methods avoid having to solve for x where there are extraneous roots.

hero (hero):

Remember, half of the class will solve for x and get it wrong.

hero (hero):

Half of the class would pick c as the answer.

OpenStudy (anonymous):

Stop showing off please..

hero (hero):

I'm not. I'm stating a fact.